OpenStudy (anonymous):

Write the expression as a single logarithm with a coefficient of 1. ln3-2ln4+ln32 I know the correct answer, I just need help with the steps needed to arrive at the answer. Thanks in advance.

OpenStudy (anonymous):

a*ln(x) = ln(x^a) ln(x)-ln(y)= ln(x/y) ln(x)+ln(y)= ln(x*y)

OpenStudy (kainui):

On your calculator do the following: ln(2)+ln(3) and see if it equals ln(6) ln(6)-ln(2) should then also be equal to ln(3) What can you figure out from this?

OpenStudy (anonymous):

Kainui, ln 2+ ln 3= Ln (2)(3), and Ln 6-Ln2= Ln (6/2) Is that right? I'm not sure why this problem is confusing me so much, I've done many others similar to this one with no trouble. Thanks for the help, hopefully I can figure it out.

OpenStudy (kainui):

Also, you're correct, but... 2ln4=ln(4^2)=ln(16) Maybe that's where you're going wrong?

OpenStudy (anonymous):

No, I got that, I think I'm just getting mixed up on my properties of logarithms. However, I just figured it out, and realized I was actually doing the problem correctly, I was just entering it into my calculator wrong. Whoops. Thanks for your help.

OpenStudy (kainui):

Cool.